a 26% aqueous solution of sulfuric acid has a density of 1.129 at 20 degrees celcius. calculate the molarity and molality
1
Expert's answer
2012-05-31T08:36:33-0400
Molarity
C(H2SO4) = n(H2SO4)/V(solution) = d(H2SO4)*w(H2SO4)/M(H2SO4) =
1.129*0.26/98 = 0.003 mol/L = 3 mM
Molality
b(H2SO4) =
n(solute)/m(solvent) = w/((1-w)*M) = 0.26/(0.74*98) = 0.0036